Show how two phase regions only need one variable to complete define the phase compositions as opposed to single phase fields which need two variables using the Gibb’s phase rule
Answer #1

From the Gibbs Phase rule,we know that the number of intesive property or number of independent state which needs to be fixed in order to fix the state of the system is given by Degrre of freedom formula (F).This is given by

F=C-P+2, where C is the number of components ,P is the number of phases.

SO if we consider a one component system i.e. C=1 for which two possible phases may exist.Then if we take two phases to exist that means P=2.So putting the value of P=2 in above equation ,we get,

F= 1-2+2= 1.So the degree of freedom gaves me the value F=1,That means I only need one intensive property to complete define the state of the system.

While if we take,the value of P=1 i.e. only one phase exist .SO calculation of degree of freedom gives us,

F=1-1+2= 2 which means we need to define two intensive property in order to fix the state of system.

In general,these 2 or one indeendent state or intensive property are Temperature and pressure of the system,and by knowing only these two properties ,we can calculate Internal energy,Enthalpy,Entropy,etc for the system.